4) A large number of voltages show a mean of 118.5V and a population standard deviation of 1.20V. Determine the percentage of data that falls between 116 and 120 Volts.
Answer #1

Given, \mu = 118.5, \sigma = 1.20

P( X < x) = p( Z < x - \mu / \sigma)

We have to find P( 116 < X < 120) = ?

P( 116 < X < 120) = P( X < 120) - P( X < 116)

= P( Z < 120 - 118.5 / 1.2) - P( Z < 116 - 118.5 / 1.2)

= P( Z < 1.25) - P( Z < -2.0833)

= 0.8944 - ( 1 - 0.9814)

= 0.8757

= 87.57%

Percentage of data falls between 116 and 120 = 87.57%